OpenStudy (anonymous):

Which of the following is the slope between the points (2, 3) and (-3, 0)?

OpenStudy (anonymous):

slope is 3/5

OpenStudy (anonymous):

an you show me how you got that last year my teacher was not an good one and I never learned

jigglypuff314 (jigglypuff314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] when given points (x1, y1) and (x2, y2) so plug in x1=2 y1=3 x2=-3 y2=0

OpenStudy (anonymous):

SO how do you get 3/5

jigglypuff314 (jigglypuff314):

when you plug it in you get 0 - 3 -3 ------- = ----- then the negatives cancel out :) (-3)-(2) -5

OpenStudy (anonymous):

o :| thank you @jigglypuff314 @erika_riera

jigglypuff314 (jigglypuff314):

glad we could help :)

OpenStudy (anonymous):

your welcome!!!:)